OpenStudy (anonymous):

if log(subscript:b)x = 0.34 and log(subscript:b)z = 0.85, evaluate the following expression: log(subscript:b)((sqrt(x)/(cubedsqrt(z))

OpenStudy (anonymous):

really need the answer please !

OpenStudy (across):

You have:\[\log_{b}(x)=0.34,\]\[\log_{b}(z)=0.85,\]\[\log_{b}(\frac{\sqrt{x}}{\sqrt[3]{z}}).\]

OpenStudy (anonymous):

\[log_{b}\frac{\sqrt{x}}{\sqrt[3]{Z}}\] use the logarithm rules to break them up into log x and log z, which you know the values of. \[log_{b}\frac{x}{y}=log_{b}x-log_{b}y\] \[log_{b}x^{a}=alog_{b}x\]

OpenStudy (anonymous):

log identities:\[\log a + \log b = \log (ab)\]\[\log a - \log b = \log (a/b)\]\[a \log b = \log b^{a}\]

OpenStudy (anonymous):

also useful: \[\sqrt[a]{x}=x^{\frac{1}{a}}\]

OpenStudy (anonymous):

is the answer -0.3642??

OpenStudy (anonymous):

That's not what I got. What expression did you eventually simplify it to?

OpenStudy (anonymous):

yeah i don't think that's write which is why i'm asking. all i did was: 0.34^(1/2) - (0.85)^(1/3)

OpenStudy (anonymous):

You're applying my 3rd formula wrong. Maybe I typed it ambiguously. a * log(b) = log(b^a)

OpenStudy (anonymous):

right*

OpenStudy (anonymous):

So how did you apply that to\[\log(x^\frac{1}{2})\]

OpenStudy (anonymous):

you are treating them like \[(log_{b}x)^{1/2}\] instead of what they actually are: \[log_{b}(x^{1/2})\]

OpenStudy (anonymous):

oh so would it be log(0.34)^(1/2) - log(0.85)^(1/3)???

OpenStudy (anonymous):

No. You need to do the manipulation before trying to plug in the actual values. Did you get to a point where you had a term like\[\log(x^{1/2})\]

OpenStudy (anonymous):

um yes..

OpenStudy (anonymous):

Use dmacine's 3rd formula rule thing before you use the actual values of log x and log y.

OpenStudy (anonymous):

And after you applied the identity log(b^a) = a log(b) what did you get?

OpenStudy (anonymous):

You can use it to turn the exponent into a multiply. You just move that exponent out front. So now apply that operation to\[\log(x^{1/2})\]